OpenStudy (anonymous):

Help Please Show all your work for finding the solution to the system of equations using subtitution? 2x+2y=38 y=x+3

OpenStudy (anonymous):

I dont want someone to do all the work just a walk through/step by step on how to get the answer.

OpenStudy (perl):

Substitution step : $$ \Large 2x + 2(x+3) = 38 $$

OpenStudy (anonymous):

so it would be come 2x+2(8)=38?

OpenStudy (perl):

$$ \Large{ 2x + 2(x+3) = 38 \\2x +2\cdot x + 2\cdot 3 = 38 } $$